OpenStudy (anonymous):

what are the solutions of the quadratic equation? X^2-9x+20=0

OpenStudy (shamil98):

(x-5)(x-4) = 0 -5 x - 4 = 20 and -5x - 4x = -9 ^_^

OpenStudy (anonymous):

20 can be written as 4*5 now if we write the equation as X^2-4X-5X+20=0 (X-5)(X-4)=0 x=5,4

OpenStudy (haseeb96):

x^2 -4x -5x +20 =0 x(x-4) -5 (x -4)=0 (x-5)(x-4)=0 either x-5=0 or x-4=0 x=5 or x=4 if you want to check the answer either it is correct or wrong put the value which you get in the given equation for instance x=5 since, x^2 -9x +20=0 5^2 -9(5)+20=0 25 -45 +20 =0 45-45=0 0=0 hence proved like this you can check the answer......